### v1

@@ III-B Robustness
Existing research primarily emphasizes LLMs robustness against malicious inputs. However, we argue that achieving robustness should also ensure instructions-following ability without exhibiting exaggerated safety, as LLMs should balance safety with helpful. In the extreme, a model that indiscriminately refuses any prompt — safe or unsafe — would be perfectly harmless but completely useless [49]. To evaluate such exaggerated safety, Röttger et al. [49] design tests using both safe prompts (which should not trigger refusals) and unsafe prompts (which should). They find that Llama2 demonstrated significant exaggerated safety; although adding guardrail prompt decrease unsafe behaviors, it simultaneously exaggerated safety. This phenomenon of exaggerated safety, which the study attributes to lexical overfitting, suggests the model may be overly sensitive to certain words or phrases. Additionally, safety-related behavior could be changed by system prompts added at inference time. An et al. [58] introduce a white-box method for automatically generating pseudo-harmful prompts to induce false refusals in LLMs. Their findings revealed that many jailbreak defenses noticeably increase the false refusal rate, undermining usability and necessitating a trade-off between minimizing such refusals and enhancing safety against jailbreaks. Similarly, Varshney et al. [48] evaluates LLMs with both safe and unsafe prompts, training classifiers to evaluate both the safety of responses and whether the model abstain or answered. They observed that “self-checking” methods improved safety but induced extremely over-defensive. Consequently, maintaining instruction-following robustness while ensuring LLM safety is an area where much work remains to be done.
@@ III-C Ethics
Ethics encompasses the moral principles and values embedded in the text generated by LLMs, ensuring that the output adheres to socially accepted moral standards and legal frameworks. Ethical content must exclude elements that are unethical, inappropriate, or misleading. Specifically, this involves avoiding language that is discriminatory, biased, or offensive, as such expressions contravene the principles of social justice [107]. A growing body of research has proposed diverse methods for evaluating the ethical compliance of content generated by LLMs, which are summarized below:
TECHHAZARDQA: [14] is a dataset created to address the security vulnerabilities of existing LLMs when generating instructional content. While prior research has primarily focused on the safety of general text-based responses, it has often overlooked the elevated risks posed by instructional outputs, such as pseudocode or code snippets. The TECHHAZARDQA dataset consists of 1,850 complex queries spanning seven technical domains, each of which can be answered in either natural language or pseudocode. The authors systematically evaluated the models’ performance across multiple response formats under various testing scenarios, including zero-shot, chain-of-thought (CoT), and few-shot learning [109, 108, 110]. To further investigate the impact of model editing on content generation, they employed the ROME model editing technique [111] to determine whether edited models were more prone to generating harmful content. The experimental results indicate that the likelihood of harmful content generation is significantly higher when models produce pseudocode compared to natural language responses. This probability further increases following model editing. These findings highlight that the generation of instructional content presents greater safety challenges than general text generation. Although model editing enhances the flexibility of LLMs, it also raises the risk of producing harmful content. The authors underscore the importance of ensuring both safety and ethical compliance when LLMs are applied to generate technical and instructional content in future developments.
LOCALVALUEBENCH: [15] is a framework designed to evaluate the value alignment and ethical performance of large language models (LLMs) across diverse cultural and legal contexts. It addresses the limitations of existing benchmarks, which are often shaped by the cultural backgrounds of their creators. This study focuses specifically on assessing the value alignment of LLMs within the Australian context, offering a demonstration framework to guide global regulators in developing localized evaluation systems.The research employs a three-tiered inquiry method to evaluate the ethical reasoning capabilities of the models: neutral questions establish baseline responses, debated questions explore how the models handle complex ethical scenarios, and misleading questions assess their behavior in extreme situations. The evaluation spans six key domains, including Capital Punishment, Weapons, and Gay Marriage, among others. To ensure experimental reliability, each response generated by the models was independently assessed by three evaluators.The results reveal distinct performance patterns among three prominent models: GPT-4, Gemini 1.5 Pro, and Claude 3 Sonet. Gemini 1.5 Pro performed well on topics such as Gay Marriage and Refugees but declined to respond to questions concerning Capital Punishment. Claude 3 Sonet demonstrated low performance on Weapons and Compulsory Voting but showed overall consistency across domains. GPT-4 underperformed on Gay Marriage, highlighting its limitations in addressing complex ethical scenarios.The analysis indicates that the cultural context embedded in the training data significantly influences the performance of each model. Furthermore, the subjective bias observed in evaluators’ ratings underscores the need for standardized evaluation frameworks to assess the ethical alignment and value sensitivity of LLMs across diverse cultural environments.
TRUSTGPT: [16] is a comprehensive framework introduced to systematically assess the ethical and social responsibility performance of large language mod(LLMs) in three key areas: toxicity, bias, and value alignment. This framework addresses gaps in existing benchmarks by providing a nuanced evaluation of these aspects through three core dimensions: toxicity analysis, bias detection, and value alignment. TRUSTGPT also measures the rate at which models choose to refuse to answer (RtA), further assessing their decision-making in morally complex scenarios. The evaluation covered eight prominent LLMs(e.g., ChatGPT, LLaMA, and Vicuna). Findings from toxicity analysis revealed that FastChat exhibited the highest toxicity levels, with substantial variation across models. Bias detection results indicated that ChatGPT demonstrated more pronounced racial and religious biases, although most models managed gender bias relatively well. Value alignment was assessed through active value alignment (AVA) and passive value alignment (PVA) tasks; ChatGPT achieved the highest soft accuracy in AVA, while other models showed significant discrepancies between soft and hard accuracy. In PVA tasks, many models struggled with conflicting norms, suggesting the need for further development in ethical alignment capabilities. The study underscores persistent challenges facing LLMs in ethics and social responsibility, particularly their tendency to generate harmful content or exhibit bias in specific scenarios. The authors emphasize the need to incorporate advanced methods, such as reinforcement learning from human feedback (RLHF), into the training process to enhance the ethical performance and alignment of LLMs.
@@
MoralBench: [20] is an innovative benchmarking framework constructed to systematically assess the moral reasoning capabilities of LLMs and their alignment with human ethical norms. While existing research has primarily focused on bias and safety, a comprehensive framework for evaluating moral judgment in LLMs has been lacking. MoralBench addresses this gap by employing the six core moral values outlined by Moral Foundations Theory [112], which include Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, and Liberty/Oppression. The evaluation framework consists of two components: Binary Moral Evaluation and Comparative Moral Evaluation. In the Binary Moral Evaluation, models are tasked with determining whether they "agree" or "disagree" with a set of moral statements, with scores based on how closely their responses align with human judgments. The Comparative Moral Evaluation requires models to select the more ethical option between two moral statements, with scoring reflecting the degree of alignment with human preferences. The experimental findings indicate that, although some models perform well on straightforward binary tasks, they encounter significant challenges with more complex comparative tasks. This suggests that models may rely more on recognizing patterns from training data rather than demonstrating a genuine understanding of ethical principles. Furthermore, the study uncovers biases within certain moral dimensions, underscoring the need for further refinement to mitigate potential risks. MoralBench offers a valuable tool for evaluating and advancing LLMs, enabling researchers to identify deficiencies in moral reasoning and informing the development of models with improved ethical performance.
@@
CVALUES: [46] addresses the limitations of existing evaluation frameworks in assessing the ethical alignment and value sensitivity of LLMs. The framework is structured around two critical dimensions: safety and responsibility. The safety dimension requires models to avoid generating harmful, biased, or unsafe content, while also ensuring their ability to manage sensitive topics, criminal advice, and privacy violations effectively. In contrast, the responsibility dimension expects models to provide positive guidance, demonstrate empathy, and account for the social and user impact of their outputs. The evaluation process integrates human assessments with automated multiple-choice evaluations, ensuring the reliability and comprehensiveness of the results. The experimental findings reveal that certain models, such as Ziya-LLaMA, offered inappropriate recommendations in response to illegal or unethical prompts, exposing vulnerabilities in their ability to handle complex scenarios safely. Furthermore, some models performed inadequately in terms of responsibility. This shortfall was particularly evident in the legal domain, where models were prone to being misled due to limited reasoning capabilities, resulting in outputs misaligned with ethical standards. Similarly, in the social sciences domain, models often failed to exhibit empathy or produce human-like responses, leading to lower responsibility scores. The study highlights that while Chinese LLMs have made significant strides in safety, their responsibility-related capabilities still require substantial improvement. CVALUES underscores the complementary nature of human and automated evaluations, advocating for their combined application to thoroughly assess models’ ethical and social responsibility performance. This approach aims to guide the development of future LLMs that are both safer and more responsible.
In addition to the studies discussed above, several other works have made important contributions to the evaluation of LLMs’ ethical performance. For instance, Sun et al. [17] developed a safety evaluation benchmark encompassing eight typical safety scenarios, including Ethics and Morality, along with six types of adversarial instruction attacks. The study introduced the SafetyPrompts benchmark to identify and assess ethical issues in LLMs models. The findings highlight that adversarial instruction attacks represent a major safety challenge for LLMs, with models exhibiting significant vulnerabilities in these contexts. The authors emphasize the need for more comprehensive safety evaluation tools and benchmarks, urging researchers to collaborate in building robust AI safety frameworks. Similarly, Zhang et al. [105] proposed a specialized framework to assess LLMs within the legal domain, focusing on ensuring both professional competence and ethical reliability in legal tasks. The results indicate that existing LLMs require further optimization when handling complex legal scenarios, particularly in terms of ethical reasoning. The authors advocate for rigorous, multi-dimensional evaluations prior to the deployment of LLMs in legal practice, emphasizing the importance of continuous optimization to reduce bias and enhance resilience against adversarial inputs. In another study, Hendrycks et al. [106] developed the ETHICS dataset to assess the moral reasoning abilities of LLMs, aiming to help align artificial intelligence more closely with shared human values. While the research shows that some models perform well on specific tasks, it also reveals substantial limitations in handling more complex ethical scenarios. The study further suggests that models’ understanding and judgment are highly sensitive to the framing of situations in text, exposing the challenges that current models face in ethical reasoning.
@@ III-D Bias and Fairness
How Are LLMs Mitigating Stereotyping Harms: [93] suggests drawing lessons from search engine technologies to mitigate the negative impact of stereotyping in LLMs-generated outputs. The researchers combined insights from natural language processing (NLP) and search engine research to develop an innovative evaluation framework. This framework evaluates LLMs handling of stereotyping associated with different social groups through autocomplete-style prompts. The authors evaluate output across four dimensions: refusal, toxicity, sentiment, and regard. The findings indicate that while the introduction of system prompts helps reduce stereotypical outputs, LLMs still struggle with certain toxic content, particularly in responses to prompts related to peoples/ethnicities and sexual orientation. Additionally, mentions of intersectional identities tend to trigger disproportionately stereotyped responses. The study further explores the implications of these findings for the convergence of LLMs and search engines, offering insights into policy development aimed at reducing stereotypes. It emphasizes the shared responsibility of model developers, scholars, NLP practitioners, and policymakers to mitigate stereotype-related harm. The paper also calls for greater awareness in areas such as training data management, leader board design and usage, and social impact measurement.
@@
Are Large Language Models Really Bias-Free: [13] investigates whether LLMs are truly bias-free. The research team conduct a comprehensive evaluation of multiple popular LLMs through two steps: an initial evaluation using standard prompts and an adversarial analysis employing jailbreak prompts. In the first step, the authors design a sentence completion task where the model is required to choose one option from two: a stereotype or an counterstereotype to complete the provided sentence. Stereotypes deemed safe in this initial evaluation will progress to adversarial analysis in the second step, where the authors employed five jailbreak techniques—role-playing, machine translation, obfuscation, prompt injection and reward incentive—aimed at bypassing language model safety filters to elicit biased responses. The evaluation measures the performance of LLMs of various sizes based on robustness, fairness, and safety. The experimental results indicate that certain categories of bias, such as those related to sexual orientation and disability, are more effectively protected by model safety measures, while biases related to gender and age are less mitigated. Compared to smaller models, medium to large models exhibit stronger performance. However, despite having 1750 billion parameters, GPT-3.5 Turbo falls below the safety threshold and displays a high degree of stereotypical responses. In addition, the adversarial analysis highlights the powerful influence of role-playing techniques, with GPT-3.5 Turbo showing vulnerability to all four attacks. Importantly, no model is entirely safe, as each is relatively susceptible to at least one jailbreak attack. Therefore, a future layered defense approach that integrates multiple safeguards may be necessary.
Is ChatGPT Fair for Recommendation: [23] explores the fairness of LLMs in recommendation systems, specifically evaluating the fairness of recommendation via LLM (RecLLM). Due to the unique nature of ReLLM, traditional fairness measurement methods based on scores and fixed datasets struggle to meet the necessary requirements. Consequently, the researchers propose a new benchmark called FaiRLLM, which includes carefully designed evaluation metrics and datasets covering eight sensitive attributes across two recommendation contexts: music and movies. The core idea of FaiRLLM is to measure the similarity between the recommendation results of neutral instructions that do not include sensitive attributes and sensitive instructions that disclose such attributes. This approach assesses fairness by analyzing the differences in similarity among various sensitive attribute values (for example, in the context of race, comparing African American, Black, White, and Asian individuals). The researchers evaluate ChatGPT using FaiRLLM and find that ChatGPT still exhibits unfairness in generating recommendations related to certain sensitive attributes. Moreover, ChatGPT’s responses align with existing societal biases regarding disadvantaged groups associated with different sensitive attributes. In addition, the authors also revealed the lack of robustness of ChatGPT to unfairness through the influence of sensitive attribute typos and language. These findings underscore that even advanced LLMs may not completely avoid social biases in recommendation tasks, emphasizing the need for more nuanced considerations and improvements regarding fairness in real-world applications.
Do Large Language Models Rank Fairly: Despite the extensive research on the efficiency and accuracy of LLMs in ranking tasks, studies on their fairness are lacking. Wang et al. [113] evaluates the fairness of LLMs as text rankers from the perspectives of users and items. Specifically, the authors focus on binary protected attributes (gender and geographic location) using the TREC Fair Ranking Track dataset and conduct two types of evaluations: list evaluation and pairwise evaluation. The listwise evaluation measures how LLMs integrate underrepresented groups into rankings from both the query-side and item-side, while the pairwise evaluation provides LLMs with two items—one from a protected group and another from a non-protected group—to compare relevance or irrelevance. The results reveal that both neural rankers and LLMs exhibit a preference for queries associated with female and European. Interestingly, Mistral-7b shows a marked bias toward male items in relevant pairs, which contrasts sharply with the behavior of other models, raising questions about the decision-making processes of these models. Overall, LLMs tend to exhibit more subtle and profound biases favoring certain protected groups.
@@
CEB: [114] proposes a unified evaluation framework to comprehensively analyze biases in LLMs. Most existing research either focuses on specific types of biases or employs incompatibility evaluation metrics, making comparisons across different datasets and LLMs challenging. To address this, the authors introduce a Compositional Evaluation Benchmark (CEB). CEB adopts a compositional taxonomy approach to describe each dataset across three dimensions: bias type, social group, and task. The bias type includes stereotyping and toxicity, social group encompass age, gender, race, and religion, while task includes direct evaluation (recognition and selection) and indirect evaluation (continuation, conversation, and classification). To facilitate compatible evaluation, the researchers establish corresponding evaluation metrics for each task, including Micro-F1 score, GPT-4-as-evaluator, Perspective API[102], Demographic Parity (DP), Equalized Odds (EO), and Unfairness Scores[115]. The experimental results reveal that GPT models achieve the best performance on direct evaluation tasks; The Refuse to Answer (RtA) rate varies across LLMs and different settings, with larger LLMs typically obtaining an RtA rate close to zero, while smaller models like Llama2 and Llama3 exhibit significantly higher RtA rates; LLMs are adept at identifying toxic content in inputs but struggle to recognize stereotypical content; Furthermore, LLMs demonstrate a high RtA rate for racial and religious social groups, indicating a heightened sensitivity; In contrast, GPT models do not stand out in terms of performance on stereotype bias types, performing comparably to smaller LLMs.
@@
In addition, there are many researchers evaluating LLMs and addressing bias and fairness issues when LLMs are used for other tasks. Huang et al. [25] introduces a Chinese bias benchmark dataset, developed through human-AI collaboration, aimed at measuring biases in Chinese LLMs. The evaluation results indicate that the bias scores for the evaluated models are significantly higher in categories such as educational qualification, disease, disability, and physical appearance compared to religion and sexual orientation. Cui et al. [4] evaluates the fairness of LLMs through a carefully designed new benchmark that utilizes the coefficient of variation as an evaluation metric across four dimensions: identity preference, credit, criminal, and health. The results show that racial groups receive the highest level of fairness. Azeem et al. [22] evaluates several highly-rated LLMs based on Human-Robot Interaction (HRI) discrimination and safety criteria, highlighting the lack of robustness in LLMs when encountering people across various protected identity characteristics (e.g., race, gender, disability status, nationality, religion, and their intersections). Zhou et al. [6] employs red teaming methods to evaluate ChatGPT, revealing that although ChatGPT performs better than other models, it is highly likely to produce biased programs with biased induction. Given the widespread use of LLMs as evaluators, Wang et al. [116] introduces a metric for conflict rate to quantitatively evaluate the model’s sensitivity to the position of responses. This novel approach indicates the presence of positional bias in LLMs, suggesting that they tend to favor responses from specific positions. For instance, GPT-4 is inclined to support responses from the first position, while ChatGPT tends to favor responses from the second position.

### v2

@@ III-B2 Benchmarking and Standardized Frameworks
Existing research primarily emphasizes LLMs robustness against malicious inputs. However, we argue that achieving robustness should also ensure instructions-following ability without exhibiting exaggerated safety, as LLMs should balance safety with helpful. In the extreme, a model that indiscriminately refuses any prompt — safe or unsafe — would be perfectly harmless but completely useless [32]. Röttger et al. [32] design XSTest, a benchmark mixing safe and unsafe prompts, and show that Llama2 exhibits exaggerated safety by refusing benign queries. They attribute this over-defensiveness to lexical overfitting and note that adding guardrail prompts, while reducing unsafe behavior, exacerbates the problem. Additionally, safety-related behavior could be changed by system prompts added at inference time. An et al. [33] propose a white-box method for automatically generating pseudo-harmful prompts, revealing that many jailbreak defenses substantially increase false refusal rates, thereby undermining model usability. Similarly, Varshney et al. [34] develop classifiers to assess whether responses are both safe and appropriately generated, finding that self-checking methods improve safety but induce excessive conservatism. These studies collectively demonstrate that robustness cannot be narrowly defined as resistance to attacks but must also account for maintaining instruction-following capacity without compromising helpfulness.
Despite substantial progress, robustness evaluation faces several unresolved challenges. Current methods still provide limited coverage across diverse domains, languages, and modalities, leaving open questions about generalization in real-world contexts. Automated adversarial testing frameworks, though increasingly sophisticated, remain inconsistent and lack standardized protocols, complicating cross-study comparison. Moreover, defenses that reduce susceptibility to jailbreaks often increase false refusals, highlighting the persistent difficulty of balancing safety with usability.
@@ III-C Ethics
Ethics refers to the moral principles and values embedded in the outputs of large language models (LLMs), ensuring that generated text adheres to socially accepted norms and legal frameworks. Ethical compliance requires excluding content that is discriminatory, offensive, or misleading, as such expressions violate the principles of fairness and social justice [131]. Existing research on ethical evaluation of LLMs can be broadly categorized into global-purpose benchmarks and local-specific evaluations.
III-C1 Global-purpose evaluation
@@
A number of works have developed comprehensive benchmarks to assess the ethical reasoning and moral alignment of LLMs. While existing research has primarily focused on bias and other safety dimensions, MoralBench [37] systematically evaluates moral reasoning by drawing on the six foundations from Moral Foundations Theory [139], which include Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, and Liberty/Oppression. It incorporates binary and comparative tasks to measure the consistency of LLM responses with human moral judgments. In the Binary Moral Evaluation, models are tasked with determining whether they “agree” or “disagree” with a set of moral statements, with scores based on how closely their responses align with human judgments. The Comparative Moral Evaluation requires models to select the more ethical option between two moral statements, with scoring reflecting the degree of alignment with human preferences. Results show that models perform adequately on simple binary classifications but struggle with comparative evaluations, suggesting a reliance on pattern recognition from training data rather than demonstrating a genuine understanding of ethical principles.
@@
CVALUES [38] addresses the lack in existing evaluation frameworks in assessing the ethical alignment and value sensitivity of Chinese LLMs. The framework is structured around two critical dimensions: safety and responsibility. The safety dimension requires models to avoid generating harmful, biased, or unsafe content, while also ensuring their ability to manage sensitive topics, criminal advice, and privacy violations effectively. In contrast, the responsibility dimension expects models to provide positive guidance, demonstrate empathy, and account for the social and user impact of their outputs. The evaluation process integrates human assessments with automated multiple-choice evaluations, ensuring the reliability and comprehensiveness of the results. While models often succeed in avoiding harmful or unsafe content, their limitations emerge in domains such as law and social sciences, where reasoning and empathy are required. This demonstrates gaps between technical safety and broader ethical responsibility.
Complementary to these efforts, Hendrycks et al. [40] proposed the ETHICS dataset to test moral reasoning in multiple scenarios, showing that model performance is highly sensitive to textual framing and limited in handling nuanced ethical dilemmas. And Sun et al. [39] developed the SafetyPrompts benchmark, which includes Ethics and Morality as one dimension among eight safety scenarios. Their findings highlight that adversarial instructions often exacerbate ethical vulnerabilities, underscoring the need for robust defenses.
III-C2 Local-Specific Evaluation
Ethics have significantly different definitions and levels of tolerance in different scenario and cultural backgrounds. Therefore, focusing on ethical risk evaluation in specific domains and localized contexts is equally important. While prior research has primarily focused on the safety of general text-based responses, TechHazardQA: [35] is a dataset created to address the safety vulnerabilities of existing LLMs when generating instructional content, such as pseudocode or code snippets. The authors systematically evaluated the models’ performance across multiple response formats under various testing scenarios, including zero-shot, chain-of-thought (CoT), and few-shot learning [140, 141, 142]. To further investigate the impact of model editing on content generation, they employed the ROME model editing technique [143] to determine whether edited models were more prone to generating harmful content. The experimental results indicate that the likelihood of harmful content generation is significantly higher when models produce pseudocode compared to natural language responses. This probability further increases following model editing.
Additionally, LocalValueBench [36] evaluates ethical alignment under localized cultural and legal contexts. Focusing on the Australian environment, it examines model responses to neutral, debated, and misleading questions across domains such as capital punishment and same-sex marriage. Neutral questions establish baseline responses, debated questions explore how the models handle complex ethical scenarios, and misleading questions assess their behavior in extreme situations. Findings reveal significant cultural sensitivity: models differ not only in performance across topics but also in how embedded cultural biases shape outputs. The analysis indicates that the cultural context embedded in the training data significantly influences the performance of each model. Furthermore, the subjective bias observed in human evaluators’ ratings expose the difficulties of standardizing cross-cultural ethical evaluation.
Taken together, global-purpose and local-specific evaluations provide complementary perspectives: the former emphasizes broad, systematic coverage of ethical reasoning, while the latter uncovers risks specific to specific domains or cultural environments. A key challenge lies in reconciling these perspectives to construct evaluation frameworks that are both comprehensive and culturally adaptive. Moreover, the diversity of ethical standards across societies raises profound challenges. Ethical standards are not universal, and LLMs trained on globally sourced data may inadvertently reflect dominant cultural values while marginalizing others. This issue of cultural diversity suggests that benchmarks must account for heterogeneous value systems rather than enforcing a single standard of morality. Looking forward, as AI systems become more autonomous, the potential emergence of machine-internal ethical reasoning poses additional risks: LLMs might develop consistent but misaligned ethical patterns that diverge from human expectations. More research is needed to address the ethical issues exposed by increasingly intelligent AI systems.
@@ III-D1 General-purpose Evaluations
A first line of research focuses on systematically probing the inherent biases embedded in LLMs through general-purpose evaluations. Leidinger et al. [47] combine insights from natural language processing (NLP) and search engine research to develop an innovative evaluation framework. This framework evaluates LLMs handling of stereotyping associated with different social groups through autocomplete-style prompts. The authors evaluate output across four dimensions: refusal, toxicity, sentiment, and regard. The findings indicate that while the introduction of system prompts helps reduce stereotypical outputs, LLMs still struggle with prompts related to peoples/ethnicities and sexual orientation. Additionally, mentions of intersectional identities tend to trigger disproportionately stereotyped responses.
@@
Furthermore, Cantini et al. [41] conduct a comprehensive evaluation of multiple popular LLMs through two steps: an initial evaluation using standard prompts and an adversarial analysis employing jailbreak prompts. In the first step, the authors design a sentence completion task where the model is required to choose one option from two: a stereotype or an counterstereotype to complete the provided sentence. Stereotypes deemed safe in this initial evaluation will progress to adversarial analysis in the second step, where the authors employe five jailbreak methods—role-playing, machine translation, obfuscation, prompt injection and reward incentive. The evaluation measures the performance of LLMs of various sizes based on robustness, fairness, and safety. The experimental results indicate that certain categories of bias, such as those related to sexual orientation and disability, are more effectively protected by model safety measures, while biases related to gender and age are less mitigated. Importantly, no model is entirely safe, as each is relatively susceptible to at least one jailbreak attack.
@@
To address the issue of most existing research either focusing on specific types of biases or employing incompatibility evaluation metrics, Wang et al. [43] introduce a Compositional Evaluation Benchmark (CEB). CEB adopts a compositional taxonomy approach to describe each dataset across three dimensions: bias type, social group, and task. The bias type includes stereotyping and toxicity, social group encompass age, gender, race, and religion, while task includes direct evaluation (recognition and selection) and indirect evaluation (continuation, conversation, and classification). To facilitate compatible evaluation, the researchers establish corresponding evaluation metrics and methods for each task, including Micro-F1 score, GPT-4-as-a-Judge, Perspective API[91], Demographic Parity (DP), Equalized Odds (EO), and Unfairness Scores[144]. The experimental results reveal that LLMs are adept at identifying toxic content in inputs but struggle to recognize stereotypical content. Furthermore, LLMs demonstrate a high refuse to answer rate for racial and religious social groups, indicating a heightened sensitivity.
@@
Similarly, Huang et al. [44] introduces a Chinese bias benchmark dataset, developed through human-AI collaboration, aimed at measuring biases in Chinese LLMs, while Cui et al. [25] evaluates the fairness of LLMs through a benchmark that utilizes the coefficient of variation as an evaluation metric across four dimensions: identity preference, credit, criminal, and health. Azeem et al. [45] evaluates several highly-rated LLMs based on Human-Robot Interaction (HRI) discrimination and safety criteria, highlighting the lack of robustness in LLMs when encountering people across various protected identity characteristics (e.g., race, gender, disability status, nationality, religion, and their intersections).
III-D2 Domain-specific Evaluations
A second line of research investigates fairness in particular downstream applications where LLMs directly affect decision-making. In the context of recommendation systems, Zhang et al. [9] evaluate the fairness of recommendation via LLM (RecLLM). Due to the unique nature of ReLLM, traditional fairness measurement methods based on scores and fixed datasets struggle to meet the necessary requirements. Consequently, the researchers propose a new benchmark called FaiRLLM, which includes carefully designed evaluation metrics and datasets covering eight sensitive attributes across two recommendation contexts: music and movies. The core idea of FaiRLLM is to measure the similarity between the recommendation results of neutral instructions that do not include sensitive attributes and sensitive instructions that disclose such attributes. This approach assesses fairness by analyzing the differences in similarity among various sensitive attribute values (for example, in the context of race, comparing African American, Black, White, and Asian individuals). The results find that ChatGPT still exhibits unfairness in generating recommendations related to certain sensitive attributes, and it’s responses align with existing societal biases regarding disadvantaged groups associated with different sensitive attributes. In addition, the authors also revealed the lack of robustness of ChatGPT to unfairness through the influence of sensitive attribute typos and language.
Similarly, Wang et al. [42] evaluates the fairness of LLMs as text rankers from the perspectives of users and items. Specifically, the authors focus on binary protected attributes (gender and geographic location) using the TREC Fair Ranking Track dataset and conduct two types of evaluations: list evaluation and pairwise evaluation. The listwise evaluation measures how LLMs integrate underrepresented groups into rankings from both the query-side and item-side, while the pairwise evaluation provides LLMs with two items—one from a protected group and another from a non-protected group—to compare relevance or irrelevance. The results reveal that both neural rankers and LLMs exhibit a preference for queries associated with female and European. Interestingly, Mistral-7b shows a marked bias toward male items in relevant pairs, which contrasts sharply with the behavior of other models, raising questions about the decision-making processes of these models. Overall, LLMs tend to exhibit more subtle and profound biases favoring certain protected groups.
In addition, Given the widespread use of LLM-as-a-Judge [145], Wang et al. [46] introduces a metric for conflict rate to quantitatively evaluate the model’s sensitivity to the position of responses. This novel approach indicates the presence of positional bias in LLMs, suggesting that they tend to favor responses from specific positions. For instance, GPT-4 is inclined to support responses from the first position, while ChatGPT tends to favor responses from the second position.
Taken together, existing research demonstrates that LLMs exhibit systematic biases both in general-purpose evaluations and in application-specific contexts. However, several limitations persist. Current benchmarks remain limited in their cultural and linguistic coverage, restricting the generalizability of fairness assessments across diverse populations. Many methods struggle to capture subtle or intersectional forms of bias, which often manifest in implicit or adversarial scenarios. Furthermore, the absence of standardized evaluation metrics and unified frameworks hinders comparability across studies, limiting cumulative progress.